OpenStudy (anonymous):

Solve the word problem. The height of a triangle with base b and area A is given by the formula h = . What is the height of a triangle with area 26 cm2 and base 8 cm? A. 1.625 cm B. 3.25 cm C. 3.5 cm D. 6.5 cm ithink its c im not sure tho

OpenStudy (alexandervonhumboldt2):

use formula base*height*area height*8=26 height=26/8=3.25

OpenStudy (alexandervonhumboldt2):

woops i meant base*height=area

OpenStudy (anonymous):

D. The formula is A = 1/2 x b x h So h = 2 x A /b h = 2 x 26 / 8 h = 52 / 8 h = 6.5 cm

OpenStudy (anonymous):

is it c or d i thought it was c..

OpenStudy (anonymous):

The area of a triangle is given as a half multiplied by the base by the height. That means that A = 0.5 x h x b Therefore my above working is correct. The first response didn't take the half into account.